A typical sheet of metal has on the average 2 defects per 5 m^2
. What is the probability that a 10 m^2 sheet of metal will have at least 3 defects?
1
Expert's answer
2019-03-18T15:29:58-0400

Assume the Poisson distribution.


"P(X = k) = e^{-\lambda}*\lambda^k/k!"

Let X denote the number of defects in a 10 m^2 sheet of metal. Then since the unit area is a 5 m^2 sheet of metal, we have


"\lambda = 2 * 10/5 = 4"

"P(X\u2a7e3)=1-e^{-4}*(4^0\/0!+4^1\/1!+4^2\/2!)""P(X\u2a7e 3)=1-13*e^{-4}=0.7619"